PLEASE HELP!!!!!!!!!!
Consider the following hypotheses:

H0: p ≥ 0.47
HA: p < 0.47

Compute the p-value based on the following sample information. (You may find it useful to reference the appropriate table: z table or t table) (Round "z" value to 2 decimal places. Round intermediate calculations to at least 4 decimal places and final answers to 4 decimal places.)

p-value
a. x = 46; n = 118 _______
b. x = 112; n = 311 _______
c. p¯ = 0.38; n = 53 _______
d. p¯ = 0.38; n = 430 _______

THANK YOU VERY MUCH!!!!!!!!!

Answer by Boreal(15235)

p hat=46/118=0.3898
z=(0.39-0.47)/sqrt(0.47*0.53/118). use the hypothesized p and 1-p, not the sample
=-0.0802/0.0459
=-1.7473
use probability that z<=1.7412 as a one way test. That is 0.0403
b has p hat=0.3601
z=-0.1099/sqrt(.47*.53/311)
=3.8832
p<0.0001
Try doing c and d using z=(0.38-0.47)/sqrt(0.47*0.53/53) and denominator of 430 for d.
then finding the probability of z to the left of that value. That is your p-value
c will have a fairly high p-value because of the smaller sample size
d will have a small p-value